OpenStudy (anonymous):

Find value of y if the line through the points (5, y) and (3, 4) has a slope of m=-7

OpenStudy (gorv):

slope=(y-4)/(5-3) =(y-4)/2 -7=(y-4)/2

OpenStudy (gorv):

multiply both side by 2

OpenStudy (gorv):

-14=y-4 add 4 on both side now that will be the solution

OpenStudy (anonymous):

thank you!!

OpenStudy (gorv):

welcm :)